OpenStudy (anonymous):

What is the solution set of 2x(x - 1) = 3?

OpenStudy (anonymous):

expand\[2x^2-2x=3\]\[2x^2-2x-3=0\] then solve using the quadratic formula

jhonyy9 (jhonyy9):

2x*2 -2x -3 =0 2 +/- radical 28 2(1+/- radical 7) 1+/- radical 7 x_1,_2 = ------------- = -------------- = ------------- 4 4 2